# PCS Workflow for Veridical Data Science in the Age of AI

### v1

@@ 1 Introduction
Data analysis as practiced today is (implicitly or explicitly) carried out through a process we call the data science life cycle (DSLC) [1]. The DSLC is a structured investigation process rooted in the context of a specific domain problem. Each step in the DSLC requires human “judgment calls” pertaining to domain knowledge and algorithmic approaches. These decisions can vary between individuals or teams, as they often involve subjective interpretations, assumptions, or preferences. As a result, different but reasonable choices can lead to different analytical paths and outcomes, introducing uncertainty into the process [2]. Traditional statistical modeling frameworks do not consider the variability stemming from judgment calls. Under-accounting these sources of uncertainty contributes in a significant way to irreproducibility of results and conclusions [3].
@@
The Predictability-Computability-Stability (PCS) framework for veridical (truthful) data science has been developed by the second author and her collaborators throughout the last decade [4, 5, 2, 6]. PCS was motivated to meet the challenges set forth by the replication crisis in the early 2010s. In particular, Prinz et al. [7] and Begley and Ellis [8] concluded that only 11-25% of major discoveries in pre-clinical oncology research at biotech companies Amgen and Bayer could be replicated. After the PCS framework was introduced in 2020 by Yu and Kumbier [5], it so happened that groups of scientists brought renewed attention to this replication issue. Recent studies from social science and ecology showed that different teams of qualified analysts’ reasonable judgment calls could lead to different data conclusions [9, 3]. This lends compelling evidence for the need of frameworks like PCS, which systematically address and quantify the impact of such judgment calls by explicitly examining the uncertainties they introduce. In a case study included in Section 9, students in a graduate applied statistics and machine learning course (STAT 214) at UC Berkeley to clean a medical dataset based on guidance from a clinician. Students made widely varying judgment calls when handling missing and inconsistent data. As a result, the predictive performance of models on each cleaned dataset showed uncertainty comparable to that induced by bootstrapping each student’s cleaned dataset individually.
@@
PCS-guided approaches have yielded successes in producing trustworthy insights and knowledge. Dwivedi et al. [10] applied PCS principles in medicine to find stable and interpretable subgroups in randomized controlled trials. In genomics, Wang et al. [11] developed a low-signal method to find experimentally validated causal genetic drivers for a heart disease called hypertrophic cardiomyopathy (HCM). Furthermore, Tang et al. [12] developed a PCS ranking procedure to select predictive genes for cost-effective cancer diagnosis. We note that while these successes are predominantly in the biomedical sciences due to the Yu Group’s close collaborations with domain experts in medicine, the PCS framework is equally applicable across domains.
@@
In this paper, we provide an updated and streamlined PCS workflow for practitioners and researchers making data-driven conclusions. This approach builds on the work of Yu and Kumbier [5] and Yu and Barter [6] while also considering the emerging contributions and implications of GenAI. The remainder of this work is structured as follows. Section 2 describes the PCS framework and its core principles. Sections 3-8 provide step-by-step instructions for the DSLC according to these PCS principles, including tips and warnings for incorporating the use of GenAI. Section 9 presents the case study described above, which uses an example from the authors’ STAT 214 course to examine uncertainty resulting from judgment calls in data cleaning. We conclude with a brief discussion on veridical data science and opportunities for future work.
@@ 2 The PCS Framework for Veridical Data Science
PCS is an evolving conceptual, philosophical, and practical research program for conducting and documenting veridical data science and AI. It rests on three core principles of data science: Predictability, Computability, and Stability [5]. These principles represent an integration of best practices and ideas from the two cultures of machine learning and statistics [13]. Furthermore, the PCS framework includes rigorously documenting the justification for decisions made throughout the DSLC using code, visualizations, and writing. An example of PCS-adherent documentation can be found in Tang and Kenney [14]. PCS for veridical data science builds “one culture” with unified language to ensure trustworthiness and safety of data science and AI.
@@
P and C are main tenets of machine learning (ML). P is used to ensure that the DSLC adequately captures reality and accurately reflects the goal of the domain problem. We thus refer to the use of P as a reality check  [^1] Predictability is used rather than Reality since early-stage development of PCS used P specifically for supervised learning.. In the supervised learning setting, predictions are a natural way to perform such a reality check [5]. Yu and Barter [6] recently extended the reality check to unsupervised learning and the rest of the DSLC by using qualitative domain knowledge and/or cross-validating discovered structure in held-out data. C goes beyond computational considerations such as speed and memory to include data-inspired simulations, as outlined in the PCS-guided simulation protocols MERITS and simChef [15, 16].
@@
S ensures that data-driven results remain stable under reasonable perturbations to the DSLC, where “reasonable” is defined within the context of the domain problem and described in documentation. This represents a substantial contribution of PCS toward establishing trust in data conclusions. As shown in Figure 1, the DSLC is an iterative process consisting of six stages. These stages begin by transforming a domain problem into a data science problem and end with communicating results effectively to a target audience. Within each stage, S involves using context-appropriate metrics to measure the impacts of reasonable perturbations. Such stability checks should be performed as frequently as resources allow. Each stability check should subsequently be thoroughly documented. Examples of suitable perturbations might include using UK BioBank data instead of Japan BioBank data or using a different but reasonable data cleaning procedure. In the modeling phase, perturbations may involve using random forests instead of gradient boosting for prediction, or using different measures of feature importance to explain the model. One special form of stability analysis is using PCS prediction perturbation intervals to report uncertainty quantification rather than relying on a single point estimate. These intervals were formally introduced for regression in Chapter 13 of Yu and Barter [6] and for multi-class classification in recent work by Agarwal et al. [17]. In both exploratory data analysis and presenting results, stability checks may entail using different types of visualizations. If sources of instability are found, S furthermore involves improving the stability of the DSLC  [^2] Works that improve stability using problem-specific information include Basu et al. [18], Wang et al. [11], & Tang et al. [12].. Thus, S is a significant expansion of the sample to sample perturbations considered in statistics under a probabilistic generative model.
@@ 2.1 PCS Use of Generative AI in Data Science
Data scientists today are frequently turning to GenAI for support throughout each stage of the DSLC. Although GenAI systems such as large language models (LLMs) can be powerful, they also have well-documented downsides. A major concern is the tendency of LLMs to hallucinate, a phenomenon in which plausible yet false information is confidently presented as fact [19]. Thus, trustworthy use of GenAI demands adherence to PCS principles in its deployment. For example, LLMs such as Gemini or Llama may be a useful way to quickly gain a high-level understanding of unfamiliar subjects relevant to the data science task. A PCS-aligned method for mitigating hallucinations is cross-referencing different GenAI systems by giving similar prompts to multiple chatbots [20]. Another suggested stability check is repeating the prompt with small perturbations (while keeping the overall message the same) in order to determine if the answer given by an LLM is consistent. Lastly, GenAI output should always be verified with trusted external sources.
@@ 3 DSLC Stage #1: Problem Formulation and Data Collection
After these goals are clear, the next step is to find available data relevant to the task. If an insufficient amount of data exists, either additional data must be gathered or the question must be adjusted. We encourage the involvement of data scientists in the data collection process when possible. This will help ensure the data is properly formatted, thoroughly documented, and able to provide strong evidence towards a conclusion. When existing data is used, we likewise encourage data scientists to understand the steps taken in data collection and how they may impact any downstream tasks. It is essential that existing data has sufficient documentation to understand its intricacies. Further readings on problem formulation and data collection can be found in Chapter 2.2 of Yu and Barter [6].
@@ 3.1 Running Example: Clinically Important Traumatic Brain Injuries
Here we introduce a real-world example that we will return to throughout this work. Traumatic brain injury (TBI) from blunt trauma is caused by an external force which moves the brain within the skull. TBI is a major cause of death in children [21, 22]. Thus, children with clinically-important traumatic brain injury (ciTBI) requiring intervention such as neurosurgery need to be quickly identified. Computed tomography (CT) scans are the standard of care for diagnosing TBI [23]. However, CT scans expose the patient to ionizing radiation. A recent study from Smith-Bindman et al. [24] found that radiation-induced cancer is estimated to result from one in every 313 head CT scans given to children. Furthermore, head CT scans contributed the largest number of cancers in children (53%) compared to other CT categories. Since only about 5% of CT scans find injuries, there are many CT scans causing needless risk to patients [25]. Thus, identifying children with near-zero risk of ciTBI can improve patient outcomes by reducing unnecessary CT scans. This was the main goal of Kuppermann et al. [25], a nationwide collaboration among physicians and researchers across various hospitals and universities in the Pediatric Emergency Care Applied Research Network (PECARN).
@@
[…]
@@ 4 DSLC Stage #2: Data Cleaning and Exploratory Data Analysis
The overarching goal of data cleaning is to obtain the version of your data that will best help reach truthful conclusions. Achieving this goal requires familiarity with both the domain problem and the intricacies of the data. Understanding the collection process and data documentation are necessary for trustworthy data cleaning, but not sufficient for full comprehension of the data’s complexities. Thus, we suggest an approach that incorporates exploratory data analysis (EDA) as a tool for thorough data cleaning. Instead of treating data cleaning like a linear task, it should be an iterative process that evolves as insights are discovered through EDA. For example, by plotting a histogram to view the distribution of a covariate, the data scientist may notice frequent missingness, suggesting imputation may be helpful for downstream modeling. Performing a stability check by plotting a boxplot instead may reveal outliers with infeasible values, uncovering possible mistakes in data entry. With each new discovery, we apply the appropriate data cleaning adjustments and return to EDA, continuing the cycle. For details such as suggested explorations and cleaning steps, we refer the reader to Chapters 4 and 5 of Yu and Barter [6].
@@
Example. Let us again consider the PECARN ciTBI data. While Kuppermann et al. [25] approached data collection thoughtfully, there are still difficulties present in the dataset. Due to the high-stakes setting of the emergency department, many patients have missing or inconsistent entries, which are likely attributable to human error. By referencing both the data documentation and the data entry form, we can better understand the missingness present in the data. The documentation notes that many covariates have “not applicable” encoded as “92”. This serves as a valuable differentiator between unintentional and intentional missingness. For example, the dataset includes an indicator for whether trauma above the clavicles is observed. In addition, six indicators denoting the location (face, neck, etc.) of such trauma are also present. It is clear from the data entry form that if a patient does not have trauma above their clavicles, all six associated indicators should intentionally have the missing value encoding of “92”. However, some measurements may be left blank, indicating unintentional missingness potentially attributable to human error. We address uncertainty stemming from data cleaning judgment calls on the ciTBI dataset more concretely in Section 9.
[…]
@@ 5 DSLC Stage #3: Uncovering Data Structure with Unsupervised Learning
Common unsupervised learning approaches include dimensionality reduction and clustering. These techniques present a challenge, since the lack of outcome labels prevents reality checks which leverage predictions. Dimensionality reduction methods such as principal component analysis (PCA) [26] help simplify the task at hand by finding a handful of features that explain the majority of the variance in the data. Stability checks of the selected features can be performed by ensuring their robustness under different data cleaning and preprocessing decisions. If labels are available, performing a principal component regression can serve as a reality check, ensuring that the features are consistent and informative for the given problem. Similarly, subgroups found by clustering methods can be assessed for consistency across multiple algorithms, including hierarchical clustering, k-means clustering, and t-SNE, to name a few [27, 28, 29, 30]. Furthermore, hyperparameters such as the number of clusters can be chosen by their ability to produce stable groupings. For example, a common approach is to evaluate the Jaccard similarity [31] of clusters obtained from different random subsamples [32]. Consistent subgroup structure provides a valuable reality check, helping ensure associated findings are robust and reliable. PCA and clustering are described in greater detail in Chapters 6 and 7 of Yu and Barter [6].
Example. The PECARN team’s goal was to identify and characterize the group of patients with near-zero risk of ciTBI. Uncovering any natural subgroup structure amongst patients may aid in detecting such a group. In the case where no stable clusters are discovered, the information gained by the resulting exploration still helps build familiarity with our data, much like EDA. However, we note that Kuppermann et al. [25] did not use unsupervised learning, opting instead to skip this stage. While this choice is not discussed in their work, a possible explanation is that unsupervised learning techniques are difficult to apply to cases with missingness or categorical features, both of which are heavily present in the ciTBI data.
[…]

### v2

@@ 1 Introduction
Data analysis as practiced today is (implicitly or explicitly) carried out through a process we call the data science life cycle (DSLC) [box1976science]. The DSLC is a structured investigation process rooted in the context of a specific domain problem. Each step in the DSLC requires human “judgment calls” pertaining to domain knowledge and algorithmic approaches. These decisions can vary between individuals or teams, as they often involve subjective interpretations, assumptions, or preferences. As a result, different but reasonable choices can lead to different analytical paths and outcomes, introducing uncertainty into the process [uncertainty2023yu, gould2025same, breznau2022data]. Traditional statistical modeling frameworks do not consider the variability stemming from judgment calls. Under-accounting these sources of uncertainty contributes in a significant way to irreproducibility of results and conclusions [gould2025same, breznau2022data].
@@
The Predictability-Computability-Stability (PCS) framework for veridical (truthful) data science has been developed by the second author and her collaborators throughout the last decade [stability2013yu, vds2020yu, uncertainty2023yu, veridical2024yu]. PCS was motivated to meet the challenges set forth by the replication crisis in the early 2010s. In particular, prinz2011believe and begley2012replication concluded that only 11-25% of major discoveries in pre-clinical oncology research at biotech companies Amgen and Bayer could be replicated. After the PCS framework was introduced in 2020 by vds2020yu, it so happened that groups of scientists brought renewed attention to this replication issue. Recent studies from social science and ecology showed that different teams of qualified analysts’ reasonable judgment calls could lead to different data conclusions [breznau2022data, gould2025same]. This lends compelling evidence for the need of frameworks like PCS, which systematically address and quantify the impact of such judgment calls by explicitly examining the uncertainties they introduce. In a case study included in Section 9, students in a graduate applied statistics and machine learning course (STAT 214) at UC Berkeley to clean a medical dataset based on guidance from a clinician. Students made widely varying judgment calls when handling missing and inconsistent data. As a result, the predictive performance of models on each cleaned dataset showed uncertainty comparable to that induced by bootstrapping each student’s cleaned dataset individually.
@@
PCS-guided approaches have yielded successes in producing trustworthy insights and knowledge. dwivedi2020stable applied PCS principles in medicine to find stable and interpretable subgroups in randomized controlled trials. In genomics, wang2024epistasis developed a low-signal method to find experimentally validated causal genetic drivers for a heart disease called hypertrophic cardiomyopathy (HCM). Furthermore, tang2025simplified developed a PCS ranking procedure to select predictive genes for cost-effective cancer diagnosis. We note that while these successes are predominantly in the biomedical sciences due to the Yu Group’s close collaborations with domain experts in medicine, the PCS framework is equally applicable across domains.
@@
In this paper, we provide an updated and streamlined PCS workflow for practitioners and researchers making data-driven conclusions. This approach builds on the work of vds2020yu and veridical2024yu while also considering the emerging contributions and implications of GenAI. The remainder of this work is structured as follows. Section 2 describes the PCS framework and its core principles. Sections 3-8 provide step-by-step instructions for the DSLC according to these PCS principles, including tips and warnings for incorporating the use of GenAI. Section 9 presents the case study described above, which uses an example from the authors’ STAT 214 course to examine uncertainty resulting from judgment calls in data cleaning. We conclude with a brief discussion on veridical data science and opportunities for future work.
@@ 2 The PCS Framework for Veridical Data Science
PCS is an evolving conceptual, philosophical, and practical research program for conducting and documenting veridical data science and AI. It rests on three core principles of data science: Predictability, Computability, and Stability [vds2020yu]. These principles represent an integration of best practices and ideas from the two cultures of machine learning and statistics [breiman2001cultures]. Furthermore, the PCS framework includes rigorously documenting the justification for decisions made throughout the DSLC using code, visualizations, and writing. An example of PCS-adherent documentation can be found in tangvdocs. PCS for veridical data science builds “one culture” with unified language to ensure trustworthiness and safety of data science and AI.
@@
P and C are main tenets of machine learning (ML). P is used to ensure that the DSLC adequately captures reality and accurately reflects the goal of the domain problem. We thus refer to the use of P as a reality check  [^1] Predictability is used rather than Reality since early-stage development of PCS used P for supervised learning.. In the supervised learning setting, predictions are a natural way to perform such a reality check [vds2020yu]. veridical2024yu recently extended the reality check to unsupervised learning and the rest of the DSLC by using qualitative domain knowledge and/or cross-validating discovered structure in held-out data. C goes beyond computational considerations such as speed and memory to include data-inspired simulations, as outlined in the PCS-guided simulation protocols MERITS and simChef [elliott2024designing, Duncan2024simChef].
@@
S ensures that data-driven results remain stable under reasonable perturbations to the DSLC, where “reasonable” is defined within the context of the domain problem and described in documentation. This represents a substantial contribution of PCS toward establishing trust in data conclusions. As shown in Figure 1, the DSLC is an iterative process consisting of six stages. These stages begin by transforming a domain problem into a data science problem and end with communicating results effectively to a target audience. Within each stage, S involves using context-appropriate metrics to measure the impacts of reasonable perturbations. Such stability checks should be performed as frequently as resources allow. Each stability check should subsequently be thoroughly documented. Examples of suitable perturbations might include using UK BioBank data instead of Japan BioBank data or using a different but reasonable data cleaning procedure. In the modeling phase, perturbations may involve using random forests instead of gradient boosting for prediction, or using different measures of feature importance to explain the model. One special form of stability analysis is using PCS prediction perturbation intervals to report uncertainty quantification rather than relying on a single point estimate. These intervals were formally introduced for regression in Chapter 13 of veridical2024yu and for multi-class classification in recent work by agarwal2025pcsuq. In both exploratory data analysis and presenting results, stability checks may entail using different types of visualizations. If sources of instability are found, S furthermore involves improving the stability of the DSLC  [^2] Stability improvements using problem-specific information include basu2018iterative, wang2024epistasis, & tang2025simplified.. Thus, S is a significant expansion of the sample-to-sample perturbations considered in statistics under a probabilistic generative model.
@@ 2.1 PCS Use of Generative AI in Data Science
Data scientists today are frequently turning to GenAI for support throughout each stage of the DSLC. Although GenAI systems such as large language models (LLMs) can be powerful, they also have well-documented downsides. A major concern is the tendency of LLMs to hallucinate, a phenomenon in which plausible yet false information is confidently presented as fact [huang2025hallucination]. Thus, trustworthy use of GenAI demands adherence to PCS principles in its deployment. For example, LLMs such as Gemini or Llama may be a useful way to quickly gain a high-level understanding of unfamiliar subjects relevant to the data science task. A PCS-aligned method for mitigating hallucinations is cross-referencing different GenAI systems by giving similar prompts to multiple chatbots [chiang2024chatbot]. Another suggested stability check is repeating the prompt with small perturbations (while keeping the overall message the same) in order to determine if the answer given by an LLM is consistent. Lastly, GenAI output should always be verified with trusted external sources.
@@ 3 DSLC Stage #1: Problem Formulation and Data Collection
After these goals are clear, the next step is to find available data relevant to the task. If an insufficient amount of data exists, either additional data must be gathered or the question must be adjusted. We encourage the involvement of data scientists in the data collection process when possible. This will help ensure the data is properly formatted, thoroughly documented, and able to provide strong evidence towards a conclusion. When existing data is used, we likewise encourage data scientists to understand the steps taken in data collection and how they may impact any downstream tasks. It is essential that existing data has sufficient documentation to understand its intricacies. Further readings on problem formulation and data collection can be found in Chapter 2.2 of veridical2024yu.
@@ 3.1 Running Example: Clinically Important Traumatic Brain Injuries
Here we introduce a real-world example that we will return to throughout this work. Traumatic brain injury (TBI) from blunt trauma is caused by an external force which moves the brain within the skull. TBI is a major cause of death in children [langlois2006tbi, faul2010tbi]. Thus, children with clinically-important traumatic brain injury (ciTBI) requiring intervention such as neurosurgery need to be quickly identified. Computed tomography (CT) scans are the standard of care for diagnosing TBI [doezema1991ct]. However, CT scans expose the patient to ionizing radiation. A recent study from smith-bindman2025cancer found that radiation-induced cancer is estimated to result from one in every 313 head CT scans given to children. Furthermore, head CT scans contributed the largest number of cancers in children (53%) compared to other CT categories. Since only about 5% of CT scans find injuries, there are many CT scans causing needless risk to patients [kuppermann2009tbi]. Thus, identifying children with near-zero risk of ciTBI can improve patient outcomes by reducing unnecessary CT scans. This was the main goal of kuppermann2009tbi, a nationwide collaboration among physicians and researchers across various hospitals and universities in the Pediatric Emergency Care Applied Research Network (PECARN).
@@
ȷ  Generative AI Tips & Warnings  . Give identical prompts to multiple different GenAI systems (Gemini, Claude, ChatGPT, etc.) to ensure stability. Change the wording of your prompt, while keeping the main message the same. Does it change the response in a meaningful way? Check reliable external sources support the LLM’s output prior to incorporating your domain knowledge into the project or task.
Key Takeaways & Action Items Ensure the data-driven task is aligned with the goals of domain experts. Verify that existing or collected data is sufficient to answer the given question. If not, problem formulation will have to be revisited. Create thorough data documentation throughout the data collection process, or ensure your previously existing data has sufficient documentation.
@@ 4 DSLC Stage #2: Data Cleaning and Exploratory Data Analysis
The overarching goal of data cleaning is to obtain the version of your data that will best help reach truthful conclusions. Achieving this goal requires familiarity with both the domain problem and the intricacies of the data. Understanding the collection process and data documentation are necessary for trustworthy data cleaning, but not sufficient for full comprehension of the data’s complexities. Thus, we suggest an approach that incorporates exploratory data analysis (EDA) as a tool for thorough data cleaning. Instead of treating data cleaning like a linear task, it should be an iterative process that evolves as insights are discovered through EDA. For example, by plotting a histogram to view the distribution of a covariate, the data scientist may notice frequent missingness, suggesting imputation may be helpful for downstream modeling. Performing a stability check by plotting a boxplot instead may reveal outliers with infeasible values, uncovering possible mistakes in data entry. With each new discovery, we apply the appropriate data cleaning adjustments and return to EDA, continuing the cycle. For details such as suggested explorations and cleaning steps, we refer the reader to Chapters 4 and 5 of veridical2024yu.
@@
Example. Let us again consider the PECARN ciTBI data. While kuppermann2009tbi approached data collection thoughtfully, there are still difficulties present in the dataset. Due to the high-stakes setting of the emergency department, many patients have missing or inconsistent entries, which are likely attributable to human error. By referencing both the data documentation and the data entry form, we can better understand the missingness present in the data. The documentation notes that many covariates have “not applicable” encoded as “92”. This serves as a valuable differentiator between unintentional and intentional missingness. For example, the dataset includes an indicator for whether trauma above the clavicles is observed. In addition, six indicators denoting the location (face, neck, etc.) of such trauma are also present. It is clear from the data entry form that if a patient does not have trauma above their clavicles, all six associated indicators should intentionally have the missing value encoding of “92”. However, some measurements may be left blank, indicating unintentional missingness potentially attributable to human error. We address uncertainty stemming from data cleaning judgment calls on the ciTBI dataset more concretely in Section 9.
ȷ  Generative AI Tips & Warnings  . LLMs can assist in efficiently creating multiple copies of clean data by providing starter code. It is important that the analyst verifies the generated code. Coding assistants such as GitHub Copilot or Cursor may be helpful for documenting and commenting data cleaning code.
Key Takeaways & Action Items Decisions made during the cleaning process should be thoroughly documented. Data cleaning should be an ongoing process that evolves as insights are uncovered through exploratory data analysis (EDA). Use EDA to identify patterns, outliers, and inconsistencies that may require further refinement of the dataset. Be cautious of over-interpreting patterns that might not generalize. Keep multiple versions of cleaned data for use in downstream stability checks.
@@ 5 DSLC Stage #3: Uncovering Data Structure with Unsupervised Learning
Common unsupervised learning approaches include dimensionality reduction and clustering. These techniques present a challenge, since the lack of outcome labels prevents reality checks which leverage predictions. Dimensionality reduction methods such as principal component analysis (PCA) [hotelling1933analysis] help simplify the task at hand by finding a handful of features that explain the majority of the variance in the data. Stability checks of the selected features can be performed by ensuring their robustness under different data cleaning and preprocessing decisions. If labels are available, performing a principal component regression can serve as a reality check, ensuring that the features are consistent and informative for the given problem. Similarly, subgroups found by clustering methods can be assessed for consistency across multiple algorithms, including hierarchical clustering, k-means clustering, and t-SNE, to name a few [ward1963cluster, lloyd1982least, macqueen1967some, vandermaaten2008tsne]. Furthermore, hyperparameters such as the number of clusters can be chosen by their ability to produce stable groupings. For example, a common approach is to evaluate the Jaccard similarity [jaccard1901index] of clusters obtained from different random subsamples [BenHur2002stability]. Consistent subgroup structure provides a valuable reality check, helping ensure associated findings are robust and reliable. PCA and clustering are described in greater detail in Chapters 6 and 7 of veridical2024yu.
Example. The PECARN team’s goal was to identify and characterize the group of patients with near-zero risk of ciTBI. Uncovering any natural subgroup structure amongst patients may aid in detecting such a group. In the case where no stable clusters are discovered, the information gained by the resulting exploration still helps build familiarity with our data, much like EDA. However, we note that kuppermann2009tbi did not use unsupervised learning, opting instead to skip this stage. While this choice is not discussed in their work, a possible explanation is that unsupervised learning techniques are difficult to apply to cases with missingness or categorical features, both of which are heavily present in the ciTBI data.
ȷ  Generative AI Tips & Warnings  . Do not solely rely on GenAI to generate pipelines for complex techniques such as PCA or clustering. Always ensure that GenAI-written code is correct. Understanding how these methods work helps you verify that the coding assistant is using them properly.
Key Takeaways & Action Items Determine if the domain problem would benefit from a certain type of underlying data structure. Incorporate stability and predictability checks into unsupervised learning to ensure validity of findings.